Hi tutors, here's a challenging 5th grade word problem which I truly have no idea to work it out.
William and Grace have $87. Grace and Siti have $115. William has 3/7 of Siti's amount. How much money does Grace have at first?
Please help. Thank you.

Answer by stanbon(75887)

William and Grace have $87. Grace and Siti have $115. William has 3/7 of Siti's amount. How much money does Grace have at first?
------
W + G = 87
S + G = 115
W = (3/7)S
-----
I'm not sure how 5th graders solve these but
this is one way to do it.
----
Substitute for W to get 2 equations in S and G
----
(3/7)S + G = 87
S + G = 115
----------------------
Subtract and solve for "S":
(4/7)S = 28
---
S = (7/4)28
S = 7*7
Siti has has $49
----
Solve for "G":
S + G = 115
49 + G = 115
Grace has $66
------
Solve for W:
W = (3/7)S
W = (3/7)49 = 3*7 = $21
=============================
Cheers,
Stan H.
==============